OpenStudy (anonymous):

solve each equation: 2^3y+1 =sqrt 2

OpenStudy (anonymous):

Recall the definition of a logarithm. If\[a^x=b\]Then\[\log_a b = x\]Try solving using that and tell me what you get.

OpenStudy (anonymous):

is 2 raised to the power 3y

OpenStudy (anonymous):

+1 as well

OpenStudy (anonymous):

we have not done logarithms yet

OpenStudy (anonymous):

I am suppose to solve with the unknown in the eponent. Does that make a difference?

OpenStudy (anonymous):

You can't do it without logarithms...

OpenStudy (anonymous):

\[2^{3y+1} = 2^{\frac{1}{2}}\] on both sides we having common base so power should be equal \[{3y+1} ={\frac{1}{2}}\] \[2 \times {3y+1} =1\] 6y + 2 = 1 6y = -1 \[y={\frac{-1}{6}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome :)